How would you prepare 100 mL of 0.035M NaCl (aq)? 


1
Expert's answer
2021-03-22T06:39:50-0400

Proportion:

0.035 mol – 1000 mL

x mol – 1000 mL

"x = \frac{0.035 \times 100}{1000} = 0.0035 \;mol"

M(NaCl) = 58.44 g/mol

"m = n \times M = 0.0035 \times 58.44 = 0.204 \;g"

You have to add 0.204 g of solid NaCl to the 100 ml volumetric flask and add water to the mark.